# Distributed Koopman Operator Learning for Perception and Safe Navigation

### v1

@@ V Conclusion
This paper has introduced a unified data-driven framework that combines distributed Koopman operator learning with MPC for predictive navigation and dynamic obstacle avoidance. By exploiting high-dimensional sensory information, the proposed approach enables multiple computational nodes to collaboratively model and forecast the motion of surrounding dynamic obstacles without centralized data fusion. The Koopman-based predictions, processed through Gaussian mixture models and polytope approximations, are embedded as linear constraints within the MPC formulation, guaranteeing safe and collision-free operation under uncertainty. The proposed method is inherently scalable and well-suited for ITS applications, where vehicles and infrastructure sensors must cooperate to perceive and react to complex, time-varying environments. The distributed learning algorithm ensures computational efficiency, reduced communication overhead, and real-time adaptability, which are key requirements for future connected and autonomous mobility platforms. Theoretical analysis confirms convergence and consistency of the distributed Koopman estimator, while simulation results demonstrate reliable, safe, and scalable motion planning performance. Overall, this work establishes a foundation for cooperative perception and distributed control in networked transportation systems, advancing the integration of learning-based modeling and control for next-generation autonomous mobility.

### v2

@@ V Conclusion
This paper presented a unified data-driven framework that integrates distributed Koopman operator learning with MPC for predictive navigation and dynamic obstacle avoidance. High-dimensional sensory data are exploited to collaboratively learn and forecast obstacle motion without centralized data fusion. The resulting Koopman-based predictions, represented via Gaussian mixtures and polytope approximations, are incorporated as linear constraints in the MPC formulation, enabling safe and collision-free navigation under uncertainty. The proposed approach is inherently scalable and well suited for autonomous navigation in complex, time-varying environments. Distributed learning ensures computational efficiency, limited communication overhead, and real-time adaptability. Convergence of the distributed Koopman estimator is theoretically guaranteed, and simulation results validate reliable, safe, and scalable motion planning. Overall, this work advances cooperative perception and distributed control for networked autonomous systems by tightly integrating learning-based modeling with predictive control.